# GreenHyperSpectra: A multi-source hyperspectral dataset for global vegetation trait prediction

### v2

@@ 7 Conclusions and perspectives
In this study, we introduce GreenHyperSpectra, a large-scale cross-sensor and cross-ecosystem spectral dataset designed to train machine learning models for plant trait prediction from hyperspectral data. Leveraging GreenHyperSpectra as a pretraining resource, we demonstrated that models using MAE consistently outperformed all other benchmarked methods, including the fully supervised baseline, across a variety of settings. The adaptability of MAE models enables their application to multi-scale remote sensing platforms, including drone, airborne, and satellite imagery, paving the way to investigate how learned spectral features, derived from a heterogeneous spectral dataset, generalize across varying spatial resolutions. MAEs also serve as a strong foundation for advanced transfer learning architectures aimed at improving predictive performance.
While we explore default sensor configurations (VNIR+SWIR and VNIR), extending pretrained encoders to other spectral ranges remains open for future work. The MAE architecture shows promise for cross-domain adaptation across heterogeneous sensing modalities through fine-tuning strategies. We contribute towards global pretraining datasets for spectral embeddings while highlighting critical biases affecting generalization. Despite improvements, run-to-run variance reveals challenges in learning stable representations from various ecological data distributions. Future research should expand multi-domain spectral datasets across biomes and sensing conditions to enhance transferability and address geographical and ecosystem-level biases in annotated data. Nevertheless, our pretrained models from GreenHyperSpectra will remain valuable as labeled data from underrepresented regions increases. This study confirms that semi- and self-supervised methods with large-scale pretraining are essential for advancing ecosystem monitoring.
@@ Appendix A Details about the datasets
For standardized cross-instrument comparison, all reflectance spectra were resampled to a uniform wavelength grid spanning the 400-2500 nm solar-reflective range. Spectral measurements were linearly interpolated to an interval of 1 nm, resulting in 2101 bands per sample. Regions of strong atmospheric water absorption, specifically 1351–1430 nm, 1801–2050 nm, and 2451–2500 nm, were removed to minimize noise and signal loss. The remaining bands were smoothed using a Savitzky-Golay filter with a 65 nm window [86]. After these steps, 1721 spectral bands were retained for analysis, providing a high-quality input space for training and evaluation.
@@
DB1[33] proximal ASD FieldSpec Pro N/A 2151 350–2500 2007 31 Reflectance spectra desert, shrubland Link
@@
DB2[36] proximal ASD FieldSpec FR N/A 1063 352–2476 2000 792 Reflectance spectra forest, shrubland: native-dominated Hawaiian forest types Link
DB3[37] proximal ASD FieldSpec FR N/A 1075 350–2498 1995 226 Reflectance spectra shrubland Link
@@
DB4[90] proximal SpecEvo PSM3500 N/A 2151 350–2500 2013 6 Reflectance spectra Forest, Gra4Sand, Shrubland, Crops Link
@@
DB5[129] proximal SVC HR-1024i 1 m 2178 338–2516 2018 112 Reflectance spectra tundra Link
DB6[101] proximal ASD FieldSpec 3 N/A 2151 350–2500 2001 112 Reflectance spectra urban vegetation Link
DB7[35] proximal ASD FieldSpec Pro N/A 210 400–2490 715 Reflectance spectra crops, gra4Sand, forest, shrubland Link
DB8[58] proximal ASD FieldSpec 3 N/A 1075 350–2498 2001 37 Reflectance spectra urban vegetation Link
DB9[68] proximal ASD FieldSpec 4 Hi-Res NG N/A 2151 350–2500 87 Reflectance spectra Link
DB10[68] proximal Beckman 5270 N/A 480 205–2976 19 Reflectance spectra Link
DB11[41] proximal SVC HR-1024i 40 cm 995 346–2499 133 Reflectance spectra aquatic vegetation Link
DB12[108] proximal SVC HR-1024i 8 cm 994 338–2515 34 Reflectance spectra tundra Link
DB13[91] proximal SVC HR-1024i N/A 2150 350–2500 2015 44 Reflectance spectra coastal, wetland Link
DB14[84] proximal ASD N/A 2101 400–2500 2021 45 Reflectance spectra urban vegetation Link
DB15[34] proximal ADS FieldSpec Pro N/A 2151 350–2500 2005 82 Reflectance spectra shrubland, steppe Link
DB16[129] proximal SVC HR-1024i 1 m 994 338–2516 2017 1660 Reflectance spectra tundra Link
DB17[102] proximal ASD FieldSpec 3 N/A 2151 350–2500 2001 490 Reflectance spectra coastal, forest, shrubland Link
DB18[79] airborne AVIRIS Classic 17–20 m 244 365–2496 2013 341 ACSR Urban, chaparral, oak woodland, conifer forest Link
@@
DB19[68] airborne AVIRIS 17–20 m 224 365–2496 32 Link
DB20[74] airborne AisaFenix 1 m 360 400–2400 2014 22889 ACSR Forest, Ecology, Land Cover, Agriculture Link
DB21[16] spaceborne Hyperion 30 m 220 400–2500 2008 2009 25 ACSR forest Link
DB22[122] airborne NEON AOP 1 m 426 380–2510 2018 10322 ACSR Link
DB23[61] airborne Hyspex 368 417–2484 2018 9993 ACSR crops Link
DB24[122] airborne NEON AOP 1 m 426 380–2510 2019 16373 ACSR Link
DB25[116] spaceborne PRISMA 30 m 69 400–2500 2021 2155 ACSR Link
DB26[116] spaceborne PRISMA 30 m 63 400–2500 2021 10000 ACSR
DB27[45] spaceborne EnMAP 30 m 224 418–2445 2022 1890 ACSR Link
@@
DB29[84] airborne AisaFenix 1 m 420 382–2499 2021 31811 ACSR urban vegetation Link
DB30[20] airborne Aviris NG 20 m 425 380–2510 2022 911 ACSR Mediterranean ecosystem Link
@@
Table 8: Descriptive statistics for plant traits in the aggregated annotated dataset[24]. List of traits and units: leaf mass per area ($\text{g}/\text{cm}^{2}$) = Cm, leaf protein content ($\text{g}/\text{cm}^{2}$) = Cp, equivalent water thickness (cm) = Cw, leaf total chlorophyll content ($\mu\text{g}/\text{cm}^{2}$) = Cab, leaf carotenoid content ($\mu\text{g}/\text{cm}^{2}$)= Car, leaf anthocyanin content ($\mu\text{g}/\text{cm}^{2}$) = Anth, Leaf Area Index ($\text{m}^{2}/\text{m}^{2}$) = LAI and carbon-based constituents ($\text{g}/\text{cm}^{2}$)= cbc (Cm-Cp).
@@ B.1 Semi-supervised regression generative adversarial network (SR-GAN)
To enable trait prediction from unlabeled hyperspectral spectra, we adopt a semi-supervised GAN framework [71]. The generator $G$ learns to synthesize spectra that are indistinguishable from real data by optimizing $\mathcal{L}_{\text{gen}}$, while the discriminator $D$ learns both to distinguish real from fake spectra and to regress plant trait values from real labeled data optimizing $\mathcal{L}_{disc}$ . We adopt this notation for $D$ outputs: $f$ is an intermediate feature extractor and $D\circ f$ is final layer trait prediction. $\text{Dist}(\cdot,\cdot)$ denotes a distance metric (e.g., cosine or Euclidean).
@@ B.2 Model Architecture Description for RTM-AE
Table 13: Parameter configuration for the PROSAIL-PRO model [43]. This table presents the default settings and corresponding notations for parameters used in the RTM block, which simulates leaf and canopy reflectance based on the PROSPECT-PRO [63] and 4SAIL [112] models.
@@ Appendix C Resource requirements
Figure 15: Feature importance of MAE-based downstream regression as a function of fine-tuning depth. Results are shown for (top) linear probing (MAE_FR_LPMAE\_FR\_LP), (middle) fine-tuning the last block (MAE_FR_9B_FTMAE\_FR\_9B\_FT), and (bottom) full fine-tuning (MAE_FR_FTMAE\_FR\_FT). The blue lines indicate the importance scores across spectral bands, while the orange line shows a reference of a vegetation spectra.

### v3

@@ 7 Conclusions and perspectives
In this study, we introduce GreenHyperSpectra, a large-scale cross-sensor and cross-ecosystem spectral dataset designed to train machine learning models for plant trait prediction from hyperspectral data. Leveraging GreenHyperSpectra as a pretraining resource, we demonstrated that models using MAE consistently outperformed all other benchmarked methods, including the fully supervised baseline, across a variety of settings. The adaptability of MAE models enables their application to multi-scale remote sensing platforms, including drone, airborne, and satellite imagery, paving the way to investigate how learned spectral features, derived from a heterogeneous spectral dataset, generalize across varying spatial resolutions. MAEs also serve as a strong foundation for advanced transfer learning architectures aimed at improving predictive performance. While we explore default sensor configurations (VNIR+SWIR and VNIR), extending pretrained encoders to other spectral ranges remains open for future work. The MAE architecture shows promise for cross-domain adaptation across heterogeneous sensing modalities through fine-tuning strategies. We contribute towards global pretraining datasets for spectral embeddings while highlighting critical biases affecting generalization. Despite improvements, run-to-run variance reveals challenges in learning stable representations from various ecological data distributions. Future research should expand multi-domain spectral datasets across biomes and sensing conditions to enhance transferability and address geographical and ecosystem-level biases in annotated data. Nevertheless, our pretrained models from GreenHyperSpectra will remain valuable as labeled data from underrepresented regions increases. This study confirms that semi- and self-supervised methods with large-scale pretraining are essential for advancing ecosystem monitoring.
@@ Appendix A Details about the datasets
For standardized cross-instrument comparison, all reflectance spectra were resampled to a uniform wavelength grid spanning the 400-2500 nm solar-reflective range. Spectral measurements were linearly interpolated to an interval of 1 nm, resulting in 2101 bands per sample. Regions of strong atmospheric water absorption, specifically 1351–1430 nm, 1801–2050 nm, and 2451–2500 nm, were removed to minimize noise and signal loss. The remaining bands were smoothed using a Savitzky-Golay filter with a 65 nm window [savitzky1964smoothing]. After these steps, 1721 spectral bands were retained for analysis, providing a high-quality input space for training and evaluation.
@@
DB1[Dennison2018RangeCreek] proximal ASD FieldSpec Pro N/A 2151 350–2500 2007 31 Reflectance spectra desert, shrubland Link
@@
DB2[DennisonGardner2018Hawaii] proximal ASD FieldSpec FR N/A 1063 352–2476 2000 792 Reflectance spectra forest, shrubland: native-dominated Hawaiian forest types Link
DB3[DennisonRoberts2018SantaMonica] proximal ASD FieldSpec FR N/A 1075 350–2498 1995 226 Reflectance spectra shrubland Link
@@
DB4[Serbin2019UWBNL] proximal SpecEvo PSM3500 N/A 2151 350–2500 2013 6 Reflectance spectra Forest, Gra4Sand, Shrubland, Crops Link
@@
DB5[Zesati2019ABoVE] proximal SVC HR-1024i 1 m 2178 338–2516 2018 112 Reflectance spectra tundra Link
DB6[Thompson2021EMITReflectance] proximal ASD FieldSpec 3 N/A 2151 350–2500 2001 112 Reflectance spectra urban vegetation Link
DB7[Dennison2019FractionalCover] proximal ASD FieldSpec Pro N/A 210 400–2490 715 Reflectance spectra crops, gra4Sand, forest, shrubland Link
DB8[Herold2004UrbanSB] proximal ASD FieldSpec 3 N/A 1075 350–2498 2001 37 Reflectance spectra urban vegetation Link
DB9[Kokaly2017USGSSpectralLib] proximal ASD FieldSpec 4 Hi-Res NG N/A 2151 350–2500 87 Reflectance spectra Link
DB10[Kokaly2017USGSSpectralLib] proximal Beckman 5270 N/A 480 205–2976 19 Reflectance spectra Link
DB11[Fang2023AquaticVeg] proximal SVC HR-1024i 40 cm 995 346–2499 133 Reflectance spectra aquatic vegetation Link
DB12[Unger2021ArcticMoss] proximal SVC HR-1024i 8 cm 994 338–2515 34 Reflectance spectra tundra Link
DB13[Serbin2015NGEE] proximal SVC HR-1024i N/A 2150 350–2500 2015 44 Reflectance spectra coastal, wetland Link
DB14[Roupioz2023CAMCATT] proximal ASD N/A 2101 400–2500 2021 45 Reflectance spectra urban vegetation Link
DB15[Dennison2018RushValley] proximal ADS FieldSpec Pro N/A 2151 350–2500 2005 82 Reflectance spectra shrubland, steppe Link
DB16[Zesati2019ABoVE] proximal SVC HR-1024i 1 m 994 338–2516 2017 1660 Reflectance spectra tundra Link
DB17[Thompson2021EMITVeg] proximal ASD FieldSpec 3 N/A 2151 350–2500 2001 490 Reflectance spectra coastal, forest, shrubland Link
DB18[Queally2018CAVeg] airborne AVIRIS Classic 17–20 m 244 365–2496 2013 341 ACSR Urban, chaparral, oak woodland, conifer forest Link
@@
DB19[Kokaly2017USGSSpectralLib] airborne AVIRIS 17–20 m 224 365–2496 32 Link
DB20[PazKagan2015Multiscale] airborne AisaFenix 1 m 360 400–2400 2014 22889 ACSR Forest, Ecology, Land Cover, Agriculture Link
DB21[Campbell2016Mongu] spaceborne Hyperion 30 m 220 400–2500 2008 2009 25 ACSR forest Link
DB22[Weinstein2022NeonTree] airborne NEON AOP 1 m 426 380–2510 2018 10322 ACSR Link
DB23[hu2023mdas] airborne Hyspex 368 417–2484 2018 9993 ACSR crops Link
DB24[Weinstein2022NeonTree] airborne NEON AOP 1 m 426 380–2510 2019 16373 ACSR Link
DB25[Vivone2022Fusion] spaceborne PRISMA 30 m 69 400–2500 2021 2155 ACSR Link
DB26[Vivone2022Fusion] spaceborne PRISMA 30 m 63 400–2500 2021 10000 ACSR
DB27[fuchs2023hyspecnet] spaceborne EnMAP 30 m 224 418–2445 2022 1890 ACSR Link
@@
DB29[Roupioz2023CAMCATT] airborne AisaFenix 1 m 420 382–2499 2021 31811 ACSR urban vegetation Link
DB30[Chadwick2025SHIFT] airborne Aviris NG 20 m 425 380–2510 2022 911 ACSR Mediterranean ecosystem Link
@@
Table 8: Descriptive statistics for plant traits in the aggregated annotated dataset[cherif2023spectra]. List of traits and units: leaf mass per area ($\text{g}/\text{cm}^{2}$) = Cm, leaf protein content ($\text{g}/\text{cm}^{2}$) = Cp, equivalent water thickness (cm) = Cw, leaf total chlorophyll content ($\mu\text{g}/\text{cm}^{2}$) = Cab, leaf carotenoid content ($\mu\text{g}/\text{cm}^{2}$)= Car, leaf anthocyanin content ($\mu\text{g}/\text{cm}^{2}$) = Anth, Leaf Area Index ($\text{m}^{2}/\text{m}^{2}$) = LAI and carbon-based constituents ($\text{g}/\text{cm}^{2}$)= cbc (Cm-Cp).
@@ B.1 Semi-supervised regression generative adversarial network (SR-GAN)
To enable trait prediction from unlabeled hyperspectral spectra, we adopt a semi-supervised GAN framework [olmschenk2019generalizing]. The generator $G$ learns to synthesize spectra that are indistinguishable from real data by optimizing $\mathcal{L}_{\text{gen}}$, while the discriminator $D$ learns both to distinguish real from fake spectra and to regress plant trait values from real labeled data optimizing $\mathcal{L}_{disc}$ . We adopt this notation for $D$ outputs: $f$ is an intermediate feature extractor and $D\circ f$ is final layer trait prediction. $\text{Dist}(\cdot,\cdot)$ denotes a distance metric (e.g., cosine or Euclidean).
@@ B.2 Model Architecture Description for RTM-AE
Table 13: Parameter configuration for the PROSAIL-PRO model [feret2021prospect]. This table presents the default settings and corresponding notations for parameters used in the RTM block, which simulates leaf and canopy reflectance based on the PROSPECT-PRO [jacquemoud1990prospect] and 4SAIL [verhoef2007unified] models.
@@ Appendix C Resource requirements
Figure 15: Feature importance of MAE-based downstream regression. Results are shown for (top) linear probing (MAE_FR_LPMAE\_FR\_LP), (middle) fine-tuning the last block (MAE_FR_9B_FTMAE\_FR\_9B\_FT), and (bottom) full fine-tuning (MAE_FR_FTMAE\_FR\_FT). The blue lines indicate the importance scores across spectral bands, while the orange line shows a reference of a vegetation spectra.